# QoE Maximization for Multiple-UAV-Assisted Multi-Access Edge Computing: An Online Joint Optimization Approach

### v1

@@ 3 System Model and Problem Formulation
At the aerial control layer, a large rotary-wing UAV (LUAV) $u$ is deployed above the service area center to serve as a regional software-defined networking (SDN) controller. It performs the following essential functions: 1) offering wireless communication coverage for UAVs at the aerial edge layer and UDs at the ground UD layer; 2) provisioning reliable edge computing services for UDs; and 3) acting as a regional controller, on which our algorithm runs, to make real-time decisions based on the collected channel state information (CSI), as well as the state information of UDs and UAVs. At the aerial edge layer, a set of small rotary-wing UAVs (SUAVs) $\mathcal{N}=\{1,\ldots,N\}$ is equipped with MEC computing capabilities to provide flexible edge computing services to UDs at the ground user layer. At the ground UD layer, a set of UDs =$\mathcal{M}=\{1,\ldots,M\}$ moving within the considered area periodically generates computing tasks with varying requirements.
In the temporal dimension, the system operates in a discrete time slot manner with $T$ equal time slots, i.e., $t\in\mathcal{T}=\{1,\ldots,T\}$, wherein each time slot has a duration of $\Delta t$. Here, $\Delta t$ is chosen to be sufficiently small such that each time slot can be considered as quasi-static. Furthermore, within each time slot, the state information of both UDs and SUAVs, as well as the CSI, are captured and updated. The task offloading, resource allocation and UAV trajectory planning are determined by running our algorithm.
3.1 Basic Models
@@
UD Model. We consider that each UD generates one computing task per time slot [30, 34]. For UD $m\in\mathcal{M}$, the attributes of the UD at time slot $t$ can be characterized as $\mathbf{St}_{m}(t)=\left(f_{m}^{\text{UD}},\mathbf{\Phi}_{m}(t),\mathbf{q}_{m}(t)\right)$, where $f_{m}^{\text{UD}}$ denotes the local computing capability of UD $m$. The computing task generated by UD $m$ is characterized as $\mathbf{\Phi}_{m}(t)=\{D_{m}(t),\eta_{m}(t),T^{\text{max}}_{m}(t)\}$ at time slot $t$, wherein $D_{m}(t)$ represents the input data size (in bits), $\eta_{m}(t)$ denotes the computation intensity (in cycles/bit), and $T^{\text{max}}_{m}(t)$ is the maximum tolerable delay (in s). $\mathbf{q}_{m}(t)=\left[x_{m}(t),y_{m}(t)\right]$ represents the location coordinates of UD $m$ at time slot $t$.
UD Mobility Model. Similar to [35], the mobility of UDs is modeled as a Gauss-Markov mobility model, which is widely employed in cellular communication networks. Specifically, the velocity of UD $m$ at time slot $t+1$ can be updated as follows:
@@
$\small\mathbf{q}_{m}(t+1)=\mathbf{q}_{m}(t)+\mathbf{v}_{m}(t)\tau.$ (2)
@@
3.2 Communication Model
The UDs can offload tasks to aerial servers [^1] Note that an aerial server and a UAV will be used interchangeably. via ground-to-air (G2A) links. To mitigate unreliable communication caused by interference, we assume that each aerial server utilizes different frequency band resources to provide edge computing services [33]. Furthermore, the widely used orthogonal frequency-division multiple access (OFDMA) technology can be employed to serve multiple UDs simultaneously [36]. Therefore, there is no communication interference in the system. Assuming UD $m$ offloads tasks to server $s\in\mathcal{S}=\{u\}\cup\mathcal{N}$ at time slot $t$, the communication rate of the UD can be calculated using Shannon’s formula as follows:
@@ 3.2 Communication Model
The probabilistic line-of-sight (LoS) channel model is employed to model the communication between the aerial servers and UDs [37]. First, the LoS probability $\rho_{s,m}^{\text{LoS}}(t)$ between server $s$ and UD $m$ at time slot $t$ can be defined as [38]
$\small\rho_{s,m}^{\text{LoS}}(t)=\frac{1}{1+c_{1}\exp(-c_{2}(\theta_{s,m}(t)-c_{1}))},$ (5)
where $c_{1}$ and $c_{2}$ are the constants depending on the propagation environment, $\theta_{s,m}(t)=\frac{180}{\pi}\arcsin{\frac{H}{d_{s,m}(t)}}$ denotes the elevation angle, and $d_{s,m}(t)$ represents the straight-line distance between server $s$ and UD $m$. Similar to [39], the path loss between server $s$ and UD $m$ can be given by
$L_{s,m}(t)=$ $20\log_{10}\left(\frac{4\pi f_{s}d_{s,m}(t)}{c}\right)+\rho_{s,m}^{\text{LoS}}(t)\eta^{\text{LoS}}$ (6)
$+\left(1-\rho_{s,m}^{\text{LoS}}(t)\right)\eta^{\text{nLoS}},$
where $f_{s}$ represents the carrier frequency of server $s$, and $c$ is the speed of light. Moreover, $\eta^{\text{LoS}}$ and $\eta^{\text{nLoS}}$ denote the additional losses for LoS and nLoS links, respectively. Without loss of generality, the channel gain $g_{s,m}(t)$ can be calculated as
$g_{s,m}(t)=10^{-\frac{L_{s,m}}{10}}.$ (7)
3.3 Computation Model
For task $\mathbf{\Phi}_{m}(t)$ generated by UD $m$, the task can be processed either locally on the UD or remotely on the aerial servers, which is determined by the offloading decision of the UD. To be more specific, we define a variable $a_{m}(t)\in\{0\}\cup\{u\}\cup\mathcal{N}$ to represent the task offloading decision of UD $m$ at time slot $t$, where $a_{m}(t)=0$ indicates that the task is executed locally, $a_{m}(t)=u$ represents that the task is offloaded to the LUAV for execution, and $a_{m}(t)=n$ means the task is offloaded to SUAV $n$ for execution.
3.3.1 Local Computing Model
@@
Task Completion Delay. The local completion latency of the task can be calculated as
$\small T_{m}^{\mathbf{loc}}(t)=\frac{\eta_{m}(t)D_{m}(t)}{f_{m}^{\text{UD}}},$ (8)
@@
$\small E_{m}^{\mathbf{loc}}(t)=k(f_{m}^{\text{UD}})^{3}T_{m}^{\text{loc}}(t),$ (9)
@@
3.3.2 Edge Computing Model
When task $\mathbf{\Phi}_{m}(t)$ is offloaded to aerial server $s\in\mathcal{S}$ for processing (i.e., $a_{m}(t)=s$), the server allocates computing resources to perform the task and returns the processing results to the UD. Note that the latency of result feedback has been disregarded when considering the task completion latency, as for most mobile applications, the size of the processing results is typically significantly smaller than the size of the input data.
@@ 3.3.2 Edge Computing Model
$T_{s,m}^{\text{ec}}(t)=\frac{D_{m}(t)}{R_{s,m}t)}+\frac{\eta_{m}(t)D_{m}(t)}{F_{s,m}(t)},$ (10)
@@
$E_{n}^{\text{c}}(t)=\sum_{m\in\mathcal{M}}I_{\{a_{m}(t)=n\}}E_{n,m}^{\text{c}}(t),$ (13)
where $I_{\{X\}}$ is an indicator function defined such that $I_{\{X\}}=1$ when event $X$ is true, and $I_{\{X\}}=0$ otherwise. Furthermore, the trajectory planning of SUAV $n$ incurs corresponding propulsion energy consumption at time slot $t$. Similar to [41, 42], the propulsion power consumption of SUAV $n$ at time slot $t$ can be expressed as
@@
3.4 QoE Evaluation Model
In this paper, we consider that the cost of each UD at time slot $t$ consists of the task completion delay and the UD energy consumption, which reflects the QoE of the UD. The completion delay of task $\mathbf{\Phi}_{m}(t)$ can be presented as
@@ 3.4 QoE Evaluation Model
Then, the energy consumption of UD $m$ can be given as
@@
Similar to [44, 45], the cost of UD $m$ at time slot $t$ can be formulated as
@@
3.5 Problem Formulation
The objective of this work is to minimize the average cost of all UDs over time (i.e., the time-averaged UD cost), by jointly optimizing the task offloading $\mathbf{A}=\{\mathcal{A}^{t}|\mathcal{A}^{t}=\{a_{m}(t)\}_{m\in\mathcal{M}}\}_{t\in\mathcal{T}}$, computing resource allocation $\mathbf{F}=\{\mathcal{F}^{t}|\mathcal{F}^{t}=\{F_{s,m}(t)\}_{s\in\mathcal{S},m\in\mathcal{M}}\}_{t\in\mathcal{T}}$, communication resource allocation $\mathbf{W}=\{\mathcal{W}^{t}|\mathcal{W}^{t}=\{w_{s,m}(t)\}_{s\in\mathcal{S},m\in\mathcal{M}}\}_{t\in\mathcal{T}}$, and trajectory planning $\mathbf{Q}=\{\mathcal{Q}^{t}|\mathcal{Q}^{t}=\{\mathbf{q}_{n}(t)\}_{n\in\mathcal{N}}\}_{t\in\mathcal{T}}$. Therefore, the problem can be formulated as follows:
@@ 3.5 Problem Formulation
Constraint (20a) is the long-term energy consumption constraint of the SUAVs. Constraint (20b) indicates that each UD can only select one strategy as its offloading decision. Constraint (20c) means that the completion delay of edge computing should not exceed the maximum tolerance delay. Constraints (20d) and (20e) imply that the allocated computing resources should be a positive value and not exceed the total amount of computing resources owned by server $s$. Constraints (20f) and (20g) limit the allocation of communication resources.
Challenges. There are two main challenges to obtain the optimal solution of problem P. i) Future-dependent. Optimally solving problem P requires complete future information, e.g., task computing demands and locations of all UDs across all time slots. However, obtaining the future information is very challenging in the considered time-varying scenario. ii) Non-convex and NP-hard. Problem P contains both continuous variables (i.e., resource allocation $\{\mathbf{F},\mathbf{W}\}$ and trajectory planning $\mathbf{Q}$) and binary variables (i.e., task offloading decision $\mathbf{A}$) is a mixed-integer non-linear programming (MINLP) problem, which is non-convex and NP-hard [46, 47].
4 Lyapunov-Based Problem Transformation
In this section, we first explain the motivation behind utilizing the Lyapunov design for online approach. Subsequently, we elaborate on the framework of the proposed Lyapunov-based online approach.
4.1 Motivation
@@ 4.2 Problem Transformation
$+\sum_{n=1}^{N}{Q}_{n}^{\mathrm{p}}(t)(E_{n}^{\mathrm{p}}(t)-\bar{E_{n}^{\mathrm{p}}})+V\times C_{s}(t),$
@@
From the inequality $(max\{a+b-c,0\})^{2}\leq(a+b-c)^{2}\ \text{for}\ \forall a,b,c\geq 0$, we can know
$Q_{n}^{\text{c}}(t+1)^{2}\leq(Q_{n}^{\text{c}}(t)+E_{n}^{\text{c}}(t)-\bar{E_{n}^{\text{c}}})^{2},$ (26)
Then, the following inequality holds, i.e.,
$\frac{Q_{n}^{\text{c}}(t+1)^{2}-Q_{n}^{\text{c}}(t)^{2}}{2}\leq\frac{(E_{n}^{\text{c}}(t)-\bar{E_{n}^{\text{c}}})^{2}}{2}+Q_{n}^{\text{c}}(t)(E_{n}^{\text{c}}(t)-\bar{E_{n}^{\text{c}}}),$ (27)
which also applies to queue $Q_{n}^{\text{p}}(t)$. Therefore, we can obtain
$\Delta L\left(\bm{\Theta}(t)\right)\leq\mathbb{E}\left\{\sum_{n=1}^{N}\frac{\left(E_{n}^{\text{c}}(t)-\bar{E_{n}^{\text{c}}}\right)^{2}+(E_{n}^{\text{p}}(t)-\bar{E_{n}^{\text{p}}})^{2}}{2}\mid\bm{\Theta}(t)\right\}$ (28)
$+\mathbb{E}\left\{\sum_{n=1}^{N}\left[Q_{n}^{\text{c}}(t)(E_{n}^{\text{c}}(t)-\bar{E_{n}^{\text{c}}})+Q_{n}^{\text{p}}(t)(E_{n}^{\text{p}}(t)-\bar{E_{n}^{\text{p}}})\right]\mid\bm{\Theta}(t)\right\}$
$\leq W+\sum_{n=1}^{N}Q_{n}^{\text{c}}(t)\left(E_{n}^{\text{c}}(t)-\bar{E_{n}^{\text{c}}}\right)+\sum_{n=1}^{N}Q_{n}^{\text{p}}(t)\left(E_{n}^{\text{p}}(t)-\bar{E_{n}^{\text{p}}}\right),$
where $W=\frac{1}{2}\sum_{n=1}^{N}\max\left\{\left(\bar{E_{n}^{\text{c}}}\right)^{2},\left(E_{\max}^{\text{c}}-\bar{E_{n}^{\text{c}}}\right)^{2}\right\}+\frac{1}{2}\sum_{n=1}^{N}\max\left\{\left(\bar{E_{n}^{\text{p}}}\right)^{2},\left(E_{\max}^{\text{p}}-\bar{E_{n}^{\text{p}}}\right)^{2}\right\}$ is a constant, and $E_{\max}^{\text{c}}$ and $E_{\max}^{\text{p}}$ are the upper bounds of computing energy consumption and propulsion energy consumption, respectively. Therefore, we have
$D(\bm{\Theta}(t))$ $\leq W+\sum_{n=1}^{N}Q_{n}^{\text{c}}(t)(E_{n}^{\text{c}}(t)-\bar{E_{n}^{\text{c}}})$ (29)
$+\sum_{n=1}^{N}Q_{n}^{\text{p}}(t)(E_{n}^{\text{p}}(t)-\bar{E_{n}^{\text{p}}})+V\times C_{s}(t).$
@@
However, problem $\textbf{P}^{\prime}$ is still a MINLP problem, with interdependencies among the decision variables. Consequently, the computational burden associated with finding the optimal solution for $\textbf{P}^{\prime}$ may not be suitable for real-time decision making. Therefore, we propose a two-stage optimization method that achieves a sub-optimal solution within a polynomial time complexity. Additionally, for the sake of convenience in subsequent discussions, we adopt a notation where the time index is omitted [49].
5 Two-Stage Optimization method
In this section, a two-stage optimization method is proposed to solve the transformed problem $\mathbf{P}^{\prime}$. In the first stage, assuming a feasible $\mathcal{Q}$, we optimize the task offloading $\mathcal{A}$ and resource allocation $\{\mathcal{F},\mathcal{W}\}$. In the second stage, based on the obtained task offloading $\mathcal{A}^{*}$ and resource allocation $\{\mathcal{F}^{*},\mathcal{W}^{*}\}$, we optimize the positions $\mathcal{Q}^{*}$ of SUAVs.
5.1 Stage 1: Task Offloading and Resource Allocation
Assuming a feasible $\mathcal{Q}$ and removing irrelevant constant terms, $\textbf{P}^{{}^{\prime}}$ can be transformed into the subproblem P1 to decide task offloading and resource allocation, which is given as follows:

### v2

@@ III System Model and Problem Formulation
At the aerial control layer, a large rotary-wing UAV (LUAV) $u$ is deployed above the service area center to serve as a regional software-defined networking (SDN) controller. It performs the following essential functions: 1) offering wireless communication coverage for UAVs at the aerial edge layer and UDs at the ground UD layer; 2) provisioning reliable edge computing services for UDs; and 3) acting as a regional controller, on which our algorithm runs, to make real-time decisions based on the collected channel state information (CSI), as well as the state information of UDs and UAVs. At the aerial edge layer, a set of small rotary-wing UAVs (SUAVs) $\mathcal{N}=\{1,\ldots,N\}$ is equipped with MEC computing capabilities to provide flexible edge computing services to UDs at the ground UD layer. At the ground UD layer, a set of UDs =$\mathcal{M}=\{1,\ldots,M\}$ moving within the considered area periodically generates computing tasks with heterogeneous and time-varying requirements.
In the temporal dimension, the system operates in a discrete time slot manner with $T$ equal time slots, i.e., $t\in\mathcal{T}=\{1,\ldots,T\}$, wherein each time slot has a duration of $\Delta t$. Here, $\Delta t$ is chosen to be sufficiently small such that each time slot can be considered as quasi-static. Furthermore, within each time slot, the state information of both UDs and SUAVs, as well as the CSI, are captured and updated. The task offloading, resource allocation and UAV trajectory control are determined by running our algorithm.
III-A Basic Models
@@
UD Model. We consider that each UD generates one computing task per time slot [30]. For UD $m\in\mathcal{M}$, the attributes of the UD at time slot $t$ can be characterized as $\mathbf{St}_{m}(t)=\left(f_{m}^{\text{UD}},\mathbf{\Phi}_{m}(t),\mathbf{q}_{m}(t)\right)$, where $f_{m}^{\text{UD}}$ denotes the local computing capability of UD $m$. The computing task generated by UD $m$ is characterized as $\mathbf{\Phi}_{m}(t)=\{D_{m}(t),\eta_{m}(t),T^{\text{max}}_{m}(t)\}$ at time slot $t$, wherein $D_{m}(t)$ represents the input data size (in bits), $\eta_{m}(t)$ denotes the computation intensity (in cycles/bit), and $T^{\text{max}}_{m}(t)$ is the maximum tolerable delay (in s). $\mathbf{q}_{m}(t)=\left[x_{m}(t),y_{m}(t)\right]$ represents the location coordinates of UD $m$ at time slot $t$.
UD Mobility Model. Similar to [33], the mobility of UDs is modeled as a Gauss-Markov mobility model, which is widely employed in cellular communication networks. Specifically, the velocity of UD $m$ at time slot $t+1$ can be updated as
@@
$\small\mathbf{q}_{m}(t+1)=\mathbf{q}_{m}(t)+\mathbf{v}_{m}(t)\Delta t.$ (2)
@@
III-B Communication Model
The UDs can offload tasks to aerial servers [^1] Note that an aerial server and a UAV will be used interchangeably. via ground-to-air (G2A) links. To mitigate unreliable communication caused by interference, we assume that each aerial server utilizes different frequency band resources to provide edge computing services [7]. Furthermore, the widely used orthogonal frequency-division multiple access (OFDMA) technology can be employed to serve multiple UDs simultaneously [34]. Therefore, there is no communication interference in the considered system. Assuming UD $m$ offloads tasks to server $s\in\mathcal{S}=\{u\}\cup\mathcal{N}$ at time slot $t$, the corresponding channel gain and transmission rate can be calculated as follows.
III-B1 Channel Gain Model
Considering that the G2A links may experience obstruction from environmental obstacles, the channel gain between UD $m$ and aerial server $s$ at time slot $t$ is calculated by incorporating the commonly used probabilistic LoS channel model [35] with the large-scale and small-scale fadings as follows:
$g_{s,m}(t)=\rho_{s,m}^{\text{LoS}}(t)g_{s,m}^{\text{LoS}}(t)+(1-\rho_{s,m}^{\text{LoS}}(t))g_{s,m}^{\text{NLoS}}(t),$ (4)
where $\rho_{s,m}^{\text{LoS}}(t)$ denotes the probability of LoS transmission. Moreover, $g_{s,m}^{x}(t)$ $(x\in\{\text{LoS},\text{NLoS}\})$ denotes the channel gain under either LoS or non-LoS (NLoS) conditions, which is expressed as $g_{s,m}^{x}(t)=|h_{s,m}^{x}(t)|^{2}L_{s,m}^{x}(t)^{-1}$, where $h_{s,m}^{x}(t)$ and $L_{s,m}^{x}(t)$ denotes the parameters of small-scale fading and large-scale fading, respectively [36].
Specifically, for the G2A links, the widely adopted LoS probability is calculated as [37]
$\rho_{s,m}^{\text{LoS}}(t)=\frac{1}{1+c_{1}\exp(-c_{2}(\frac{180}{\pi}\arcsin{\frac{H}{d_{s,m}(t)}}-c_{1}))},$ (5)
where $c_{1}$ and $c_{2}$ are the constants depending on the environment, and $d_{s,m}(t)$ means the straight-line distance between UD $m$ and server $s$. Moreover, the small-scale fading for G2A links at time slot $t$ is modeled as a parametric-scalable and good fitting generalized fading, i.e., Nakagami-$m$ fading [38], which is given as
$h_{s,m}^{x}(t)$ $\sim f^{\mathrm{Nak}}\left(h_{s,m}^{x}(t),\theta^{x}\right)$ (6)
$=\frac{2(\theta^{x})^{\theta^{x}}}{\Gamma(\theta^{x})(\bar{p})^{\theta^{x}}}(h_{s,m}^{x}(t))^{2\theta^{x}-1}\exp\big{(}-\frac{\theta^{x}}{\bar{p}}(h_{s,m}^{x}(t))^{2}\big{)},$
where $\bar{p}$ is the average received power, $\Gamma(\cdot)$ is the Gamma function, and $\theta^{x}$ $(x\in\{\text{LoS},\text{NLoS}\})$ is the Nakagami-$m$ fading parameters for LoS/NLoS channel. Furthermore, the large-scale fading for the G2A links can be given as [39]
$L_{s,m}^{x}(t)=\frac{\left(4\pi d_{s,m}(t)f_{s}\right)^{2}}{c^{2}}\eta^{x},$ (7)
where $c$ denotes the speed of light, $f_{s}$ is the carrier frequency, and $\eta^{x}$ is the attenuation factor for LoS/NLoS channel.
III-B2 Transmission Rate Model
Based on the above analysis, the transmission rate between UD $m$ and server $s$ can be calculated using Shannon’s formula as follows:
@@ III-B2 Transmission Rate Model
III-C Computation Model
For task $\mathbf{\Phi}_{m}(t)$ generated by UD $m$, the task can be processed either locally on the UD or remotely on the aerial servers, which is determined by the offloading decision of the UD. To be more specific, we define a variable $a_{m}(t)\in\{0\}\cup\{u\}\cup\mathcal{N}$ to represent the task offloading decision of UD $m$ at time slot $t$, where $a_{m}(t)=0$ indicates that the task is executed locally, $a_{m}(t)=u$ represents that the task is offloaded to the LUAV for execution, and $a_{m}(t)=n\in\mathcal{N}$ means the task is offloaded to SUAV $n$ for execution.
III-C1 Local Computing Model
@@
Task Completion Delay. The local completion delay of the task can be calculated as
$\small T_{m}^{\text{loc}}(t)=\frac{\eta_{m}(t)D_{m}(t)}{f_{m}^{\text{UD}}},$ (9)
@@
$\small E_{m}^{\text{loc}}(t)=k(f_{m}^{\text{UD}})^{3}T_{m}^{\text{loc}}(t),$ (10)
@@
III-C2 Edge Computing Model
When task $\mathbf{\Phi}_{m}(t)$ is offloaded to an aerial server $s\in\mathcal{S}$ for processing (i.e., $a_{m}(t)=s$), the server allocates computing resources to perform the task and returns the processing results to the UD. Note that, similar to most existing studies [18, 41], the latency of result feedback can be disregarded when considering the task completion delay. This is because for many intelligent applications (such as image recognition, real-time video analytics), the size of the processing results is typically significantly smaller than the size of the input data.
@@ III-C2 Edge Computing Model
$\small T_{s,m}^{\text{ec}}(t)=\frac{D_{m}(t)}{R_{s,m}t)}+\frac{\eta_{m}(t)D_{m}(t)}{F_{s,m}(t)},$ (11)
@@
$\small E_{n}^{\text{c}}(t)=\sum_{m\in\mathcal{M}}I_{\{a_{m}(t)=n\}}E_{n,m}^{\text{c}}(t),$ (14)
where $I_{\{X\}}$ is an indicator function defined such that $I_{\{X\}}=1$ when event $X$ is true, and $I_{\{X\}}=0$ otherwise. Furthermore, the trajectory control of SUAV $n$ incurs corresponding propulsion energy consumption at time slot $t$. Similar to [43], the propulsion power consumption of SUAV $n$ at time slot $t$ can be expressed as
@@
III-D QoE Evaluation Model
Given the limited battery capacity of UDs and the latency sensitivity of computing tasks, task completion delay and UD energy consumption are important performance indicators for evaluating the QoE of UDs. Similar to [45], we combine task completion latency and UD energy consumption into a cost function to measure the QoE of UDs. Specifically, the completion delay of task $\mathbf{\Phi}_{m}(t)$ at time slot $t$ can be given as
@@ III-D QoE Evaluation Model
Then, the energy consumption of UD $m$ at time slot $t$ can be given as
@@
According to [45], the cost of UD $m$ at time slot $t$ can be formulated as
@@
III-E Problem Formulation
The objective of this work is to minimize the average cost of all UDs over time (i.e., the time-averaged UD cost), by jointly optimizing the task offloading $\mathbf{A}=\{\mathcal{A}^{t}|\mathcal{A}^{t}=\{a_{m}(t)\}_{m\in\mathcal{M}}\}_{t\in\mathcal{T}}$, computing resource allocation $\mathbf{F}=\{\mathcal{F}^{t}|\mathcal{F}^{t}=\{F_{s,m}(t)\}_{s\in\mathcal{S},m\in\mathcal{M}}\}_{t\in\mathcal{T}}$, communication resource allocation $\mathbf{W}=\{\mathcal{W}^{t}|\mathcal{W}^{t}=\{w_{s,m}(t)\}_{s\in\mathcal{S},m\in\mathcal{M}}\}_{t\in\mathcal{T}}$, and trajectory control $\mathbf{Q}=\{\mathcal{Q}^{t}|\mathcal{Q}^{t}=\{\mathbf{q}_{n}(t)\}_{n\in\mathcal{N}}\}_{t\in\mathcal{T}}$. Therefore, the problem can be formulated as follows:
@@ III-E Problem Formulation
Constraint (21a) ensures that the long-term energy consumption of each SUAV does not exceed its designated energy budget. Constraint (21b) defines the strategy space for task offloading decisions of each UD. Constraint (21c) guarantees that the task completion delay for edge computing does not exceed the maximum tolerable latency of the task. Moreover, constraints (21d) and (21e) specify that the amount of computing resources allocated by each aerial server must be positive and should not exceed its total available computing resources. In addition, constraints (21f) and (21g) require that the communication resource allocation coefficients of each aerial server be positive and remain within the available communication resources of the server. Finally, constraint (21h) limits the initial positions of SUAVs, the maximum flying distance per time slot, and the minimum safe distance between any pair of SUAVs.
Challenges. There are two main challenges to obtain the optimal solution of problem P. i) Future-dependent. Optimally solving problem P requires complete future information, e.g., task computing demands and locations of all UDs across all time slots. However, obtaining the future information is very challenging in the considered time-varying scenario. ii) Non-convex and NP-hard. Problem P contains both continuous variables (i.e., resource allocation $\{\mathbf{F},\mathbf{W}\}$ and trajectory control $\mathbf{Q}$) and binary variables (i.e., task offloading decision $\mathbf{A}$) is a mixed-integer non-linear programming (MINLP) problem, which is non-convex and NP-hard [46, 47]. To this end, we proposed OJTRTA to solve this problem, which consists of a Lyapunov-based problem transformation process and a two-stage optimization procedure. Fig. 2 shows the framework of the proposed OJTRTA.
Figure 2: The framework of OJTRTA. The original problem P, which depends on future information, is first transformed into a per-slot real-time optimization problem $\textbf{P}^{\prime}$ using the Lyapunov optimization framework. Then, the problem $\textbf{P}^{\prime}$ is decomposed into two subproblems: the task offloading and resource allocation subproblem P1, and the UAV trajectory control subproblem P2. For P1, the optimal resource allocation strategy is obtained via convex optimization, while the task offloading decision is obtained using game-theoretic methods. For P2, the UAV trajectory control strategy is obtained by reformulating the problem into a convex optimization form via the SCA technique, which is subsequently solved by convex optimization.
IV Lyapunov-Based Problem Transformation
In this section, we first explain the motivation for designing an online approach using the Lyapunov optimization framework. Then, we provide a detailed explanation of the Lyapunov-based problem transformation.
IV-A Motivation
@@ IV-B Problem Transformation
$+\sum_{n=1}^{N}{Q}_{n}^{\mathrm{p}}(t)(E_{n}^{\mathrm{p}}(t)-\bar{E_{n}^{\mathrm{p}}})+V\cdot C_{s}(t),$
@@
Refer to Section I of the supplementary material.
@@
However, problem $\textbf{P}^{\prime}$ is still a MINLP problem, with interdependencies among the decision variables. Note that for this complex problem $\textbf{P}^{\prime}$, obtaining the optimal solution may not be suitable for the considered scenario. First, although the optimal solution can ensure the maximization QoE for all UDs, it may result in unfairness among UDs. Second, obtaining the optimal solution may result in significant computational overhead, which is difficult to meet the needs of real-time decision making. To this end, we propose a two-stage optimization method that uses game theory to determine the offloading decisions of UDs to ensure fairness among UDs. Moreover, the proposed method can also achieve a high-performance solution within a low computational complexity. The proposed two-stage optimization method is described in detail in the following section. Furthermore, for the sake of convenience in subsequent discussions, we adopt a notation where the time index is omitted.
V Two-Stage Optimization method
In this section, a two-stage optimization method is proposed to solve the transformed problem $\mathbf{P}^{\prime}$. In the first stage, assuming a feasible $\mathcal{Q}$, we optimize the task offloading $\mathcal{A}$ and resource allocation $\{\mathcal{F},\mathcal{W}\}$. In the second stage, based on the obtained task offloading $\mathcal{A}^{*}$ and resource allocation $\{\mathcal{F}^{*},\mathcal{W}^{*}\}$, we optimize the positions $\mathcal{Q}^{*}$ of SUAVs. It should be noted that although we optimize the resource allocation, task offloading, and UAV trajectory control in a decoupled manner, the decision variables related to these aspects are still considered together throughout the decoupling process. This is because optimizing one decision variable is conducted under the assumption of given values for the other two variables.
V-A Stage 1: Task Offloading and Resource Allocation
Assuming a feasible $\mathcal{Q}$ and removing irrelevant constant terms, $\textbf{P}^{{}^{\prime}}$ is transformed into the subproblem P1 to decide task offloading and resource allocation, which is given as follows: